OpenStudy (anonymous):

The value of Mr. Cox’s car x years after its purchase is given by the function, V(x) = 20,000(0.85)x. Approximately, what was the value of Mr. Cox’s car 5 years after its purchase?

OpenStudy (ganpat):

V(x) = 20,000(0.85)x. substitute, x = 5 in the equation and simplify to find the value...

OpenStudy (anonymous):

well i did 20000(0.85)^5 = 8874 what do i do after that

OpenStudy (ganpat):

nopes.. multiply.. do not raise to th epower..

OpenStudy (anonymous):

what is it i multiply?

OpenStudy (ganpat):

V(x) = 20,000(0.85)x. V(5) = 20,000 (0.85) * 5 V (5) = 85000

OpenStudy (anonymous):

so the answer is V(5)=85000?

OpenStudy (ganpat):

yup.. u understood ??

OpenStudy (anonymous):

yep i sure do thanks :)

OpenStudy (ganpat):

w/c,.